OpenStudy (anonymous):

A uniform plank of length 4.0 m and weight 226 N rests horizontally on two supports, with 1.1 m of the plank hanging over the right support (see the drawing). To what distance x can a person who weighs 454 N walk on the overhanging part of the plank before it just begins to tip?

OpenStudy (sriram):

wat drawing?

OpenStudy (sriram):

conserve the torque 226*1.1=454*x x=1.1 / 2=5.5 m from the fulcrum

OpenStudy (sriram):

oops i m sorry but is there another fulcrum in the left??

OpenStudy (anonymous):

yeah- there is one on the far left then one 1.1 m from the right

OpenStudy (sriram):

i m so sorry i misread the question

OpenStudy (sriram):

k i got it there are 2 fulcrums or supports the support on the LHS ensures that the plank won,t fall or rotate in anti clockwise direction.

OpenStudy (sriram):

so for movement to start it has to rotate in clockwise direction

OpenStudy (sriram):

hence for the torque eq. (look at support on RHS) weight on its right= (1.1/4.0) *226 n at distance 1.1/2 from it weight on its left=(2.9/4.0)*226 n at distance 2.9/2 from it

OpenStudy (sriram):

let d man be at a distance x on it's RHS now write d torque eq..

OpenStudy (sriram):

(1.1/4)*226*(1.1/2) + x*454 = (2.9/4)*226*(2.9/2) solve for x if man moves any further from this the plnck will rotate

OpenStudy (sriram):

i hope u understood

OpenStudy (sriram):

good question